OpenStudy (anonymous):

-6+√x-5=-2 i moved -6 over to the other side so now it looks like: √x-5=4 now what?

OpenStudy (anonymous):

square on both sides

OpenStudy (anonymous):

*squaring

OpenStudy (anonymous):

so the square cancels out the check mark thing that i cant remember the name of?

jimthompson5910 (jim_thompson5910):

square root

OpenStudy (anonymous):

\[(\sqrt{x-5})^2=(4)^2\]

jimthompson5910 (jim_thompson5910):

\[\large -6+\sqrt{x-5}=-2\] \[\large \sqrt{x-5}=-2+6\] \[\large \sqrt{x-5}=4\] \[\large x-5=4^2\] \[\large x-5=16\] \[\large x=16+5\] \[\large x=21\]